# Fibonacci-Modulation-Induced Multiple Topological Anderson Insulators

### v1

@@ Abstract
Localization properties.—A commonly held view is that in the TAI phase, all states are localized, exhibiting Anderson localization. However, in our case, the Fibonacci modulation fundamentally alters the localization properties. Nearly all the states demonstrate multifractal characteristics for any nonzero $\lambda$, irrespective of the topological phase. Figure 3 shows the density distributions of different eigenstates in distinct topological phases in both real and momentum spaces, respectively. The momentum-space density distribution is defined as $n_{k}=1/L\sum_{i,j}e^{-ik(i-j)}(\rho^{\uparrow}_{ij}+\rho^{\downarrow}_{ij})$, where $\rho^{(\beta)}_{ij}=\langle\Psi_{n}|c^{{\dagger}}_{i\beta}c_{j\beta}|\Psi_{n}\rangle$ represents the single-particle density matrix of the $n$th eigenstate $|\Psi_{n}\rangle=\sum_{i,\beta}\psi_{i\beta}^{(n)}|i,\beta\rangle$, with the eigenvalue $E_{n}$. The real-space density distribution is given by $n_{i}=\rho_{ii}^{(\uparrow)}+\rho_{ii}^{(\downarrow)}$. In momentum space, a localized (extended) state exhibits an extended (localized) distribution. Multifractal states, however, exhibit delocalized yet nonergodic behavior in both real and momentum spaces. We select the $198$th eigenstates for $\lambda=2$ in the topologically trivial phase and $\lambda=3$ in the TAI phase with $L=610$ and $\phi=0$ under OBC, as shown in the top and bottom rows of Fig. 3, respectively. The density distributions in both spaces confirm the multifractal behavior, which persists regardless of whether the states is localized in topologically nontrivial phase or not.
@@
To characterize the localization properties of the wave function, we calculate the fractal dimension of $n$th eigenstate using $D_{n}=-\lim_{L\to\infty}\ln{(\mathrm{IPR}_{n})}/\ln{(2L)}$, with the inverse participation ratio $\mathrm{IPR}_{n}=\sum_{i,\beta}|\psi_{i\beta}^{(n)}|^{4}$. For a localized (extended) state, $D_{n}\to 0$ ($D_{n}\to 1$) in the thermodynamic limit. When $10<D_{n}<$ 1, the state is considered multifractal. Figures 4(a) and 4(b) show the fractal dimensions for different system sizes under periodic boundary condition (PBC) with $\phi=0$ for $\lambda=2$ and $\lambda=3$, respectively. For both topologically trivial and nontrivial phases, the fractal dimensions of nearly all states are system-size independent, deviating from the extremes of 00 and $1$. The scaling behavior of the mean fractal dimension $\langle D\rangle=1/(2L)\sum_{n}D_{n}$ for various $\lambda$ values, in both topologically trivial ($\lambda=2$, $3.5$, and $4.2$) and nontrivial ($\lambda=3$, $3.8$, and $4.8$) phases, is shown in Fig. 4(c). In the thermodynamic limit, $\langle D\rangle$ approaches values that deviate from both 00 and $1$ in both phases. Our results suggest that the emergence of the TAI phase does not affect the localization properties of the systme, and the states remain multifractal.
Figure 5: (Color online) (a) $\overline{C}(t)$ versus $t$ for different $\lambda$ with $N_{c}=50$ disorder realizations. (b) $\langle\overline{C}\rangle$ as a function of $\lambda$ with $N_{c}=200$ disorder realizations. (c) The evolution of $\Lambda(t)$ with $\lambda_{f}=0$, calculated for the $(L+1)$th eigenstate at different values of $\lambda_{i}$ as the initial states. (d) The evolution of $W(t)$ for different $\lambda$ with $\phi=0$. Here, $L=144$, $t_{so}=0.5$, and $M=-4$.
@@
Dynamical detection.—Experimentally, this Fibonacci-modulated SOC chain can be realized along the momentum lattice [84, 85, 86, 37, 87, 88, 89, 90, 91, 92, 93, 94] in a cold atom of Bose-Einstein condensate [83], and its topological properties and localization features can be detected by the quench dynamics. To characterize the topology of our model, we monitor the dynamical response of the system to a sudden quench. Specifically, we measure the mean chiral displacement [37, 34, 51, 95, 42] and the Loschmidt echo (LE)[97, 96, 99, 98, 100, 101]. We define the disorder-averaged expectation value of the chiral displacement operator as $\overline{C}(t)=2\langle\Psi(t)\left|\Gamma X\right|\Psi(t)\rangle$, where $\Gamma$ is the chiral operator, $X$ is the unit cell operator, and $|\Psi(t)\rangle=e^{iHt}|\Psi(0)\rangle$ is the time-evolved wave function, with the initial state $|\Psi(0)\rangle=c_{j_{0}\uparrow}^{\dagger}|0\rangle$ and $j_{0}$ being the position of the central bulk lattice site. The dynamics of $\overline{C}(t)$ generally exhibit transient, oscillatory behavior, as shown in Fig. 5(a) for different $\lambda$ with $M=-4$ and $t_{so}=0.5$. In long time limit, $\bar{C}(t)$ converges to 00 for $\lambda$ values corresponding to the topologically trivial phase and to $-1$ for those corresponding to the nontrivial phase. To eliminate the oscillation, we compute the time average $\langle\overline{C}\rangle$, which converges to $-1$ for a topologically nontrivial phase and to 00 for a trivial one. Figure 5(b) shows $\langle\overline{C}\rangle$ as the function of the modulation amplitude $\lambda$ with $M=-4$ and $t_{so}=0.5$. The value of $\langle\overline{C}\rangle$ starts at 00, then jumps to a nontrivial value three times, before eventually returning to 00. These results demonstrate that the mean chiral displacement is a sensitive probe of the multiple TAI phases in our model, effectively capturing the topological transitions driven by the modulation amplitude $\lambda$.
@@
The LE, defined as $\mathcal{L}(t)=|\langle\Psi(0)|e^{-iH(\lambda_{f})t}|\Psi(0)\rangle|^{2}$, is a powerful tool for analyzing nonequilibrium dynamics. It exhibits a series of zero points at specific time intervals when the initial Hamiltonian [$H(\lambda_{i})$] and the post-quench Hamiltonian [$H(\lambda_{f})$] describe systems that are localized in different phases. This behavior is a hallmark of dynamical quantum phase transitions (DQPTs), which have been observed in various systems[102, 103, 104, 105, 106, 107, 108]. Pioneering studies have established a connection between DQPTs and emergent topological transitions. In particular, dynamical topological order parameters can change their integer values when DQPTs occur, providing a means to dynamically track the topological changes during the quench. The appearance of zero points in the LE signals the nonanalytic behavior in the dynamical free energy $\Lambda(t)=-1/(2L)\ln\mathcal{L}(t)$. Figure 5(c) shows the evolution of $\Lambda(t)$ with $\lambda_{f}=0$, calculated for the $(L+1)$th eigenstate at different values of $\lambda_{i}$ for $M=-4$ and $t_{so}=0.5$ as the initial states. For the topologically trivial Hamiltonian $H(\lambda_{f})$, when the initial Hamiltonian $H(\lambda_{i})$ is in the topologically nontrivial phase ($\lambda_{i}=3$, $3.8$, and $4.8$), a serial of divergence points of $\Lambda(t)$–corresponding to exact zeros of LE–emerge along the time axis. In contrast, for $\lambda_{i}$ values within the topologically trivial phase ($\lambda_{i}=2$, $3.5$, and $4.2$), the nonanalytic behavior disappears, indicating that both the initial and the post-quench Hamiltonians are localized in the same topological phase. These results suggest that the LE method not only captures the occurrence of DQPTs but also provides a powerful tool to probe the emergence of multiple TAI phases in our system.
@@
To dynamically detect the localization properties of our system, we employ the mean-square displacement[34, 109, 110, 2] defined as
@@
Conclusion.—We have demonstrated that multiple TAI phases can be induced by Fibonacci modulation in a SOC chain. Unlike conventional TAI phases, where the system exhibits full localization, our results show that the system retains multifractal features, regardless of whether it is in the TAI phase or not. Crucially, the emergence of multiple TAI phases is closely tied to the fractal structure of the spectrum induced by the Fibonacci modulation. The self-similar splitting of the bands at each scale enables the formation of multiple TAI phases, resulting in a topologically nontrivial structure that evolves as the SOC amplitude is varied. This model can be experimentally realized along the momentum lattice in a cold atomic system, and its properties can be tested through dynamical evolution experiments.
Acknowledgements.

### v3

@@ Introduction
A commonly held view is that in the TAI phase, all states are localized, exhibiting Anderson localization. However, in our case, the Fibonacci modulation fundamentally alters the localization properties. For any nonzero $\lambda$, nearly all the states display multifractal characteristics, regardless of the topological phase. Figure 3 shows the density distributions of several representative eigenstates in real and momentum spaces for the case of $\kappa=1$ with different $\lambda$ where the system is localized in both topologically trivial ($\lambda=2$, $3.5$, and $4.2$) and nontrivial ($\lambda=3$, $3.8$, and $4.8$) phases. Here, we select $n=198$, $345$, and $528$ as examples, where the eigenvalues $E_{n}$ are ordered in ascending order. The momentum-space density distribution Wang2020 is defined as $n_{k}=1/L\sum_{i,j}e^{-ik(i-j)}(\rho^{\uparrow}_{ij}+\rho^{\downarrow}_{ij})$, where $\rho^{(\beta)}_{ij}=\langle\Psi_{n}|c^{{\dagger}}_{i\beta}c_{j\beta}|\Psi_{n}\rangle$ represents the single-particle density matrix of the $n$th eigenstate, $|\Psi_{n}\rangle=\sum_{i,\beta}\psi_{i\beta}^{(n)}|i,\beta\rangle$, with the eigenvalue $E_{n}$. The real-space density distribution is given by $n_{i}=\rho_{ii}^{(\uparrow)}+\rho_{ii}^{(\downarrow)}$. In momentum space, a localized (extended) state exhibits an extended (localized) distribution Liutong2025 ; DaiQ2023 . Multifractal states, however, exhibit delocalized yet nonergodic behavior in both real and momentum spaces. The density distributions in both spaces confirm the multifractal nature of the states, which persists irrespective of whether the states are in a topologically nontrivial phase or not.
Figure 4: Fractal dimensions in both real and momentum spaces. Fractal dimensions for different system sizes under PBCs in both real (a1)-(a6) and momentum (b1)-(b6) spaces with the spin-flip hopping strength $t_{so}=0.5$, the uniform Zeeman potential M=4M=-4, $\kappa=1$, and $\phi=0$. From left to right, the columns correspond to different quasiperiodic modulation amplitudes $\lambda=2.0$, $3.0$, $3.5$, $3.8$, $4.2$, and $4.8$, where the system is localized in different topological regimes.
Figure 5: The scaling of the mean fractal dimensions in both real and momentum spaces. The scaling of the mean fractal dimensions in both (a) real and (b) momentum spaces for various values of the quasiperiodic modulation amplitude $\lambda$. The dashed lines correspond to the finite-size extrapolation of $\langle D_{r}\rangle$ ($\langle D_{k}\rangle$) as a function of $1/\ln(2L)$. Here, the spin-flip hopping strength $t_{so}=0.5$, the uniform Zeeman potential M=4M=-4, $\kappa=1$, and $\phi=0$.
@@
By contrasting the fractal dimensions for each eigenstate at different system sizes in both real and momentum spaces, we can clearly characterize the localization properties of our system across distinct topological phases as the modulation amplitude increases. This analysis is illustrated in Fig. 4 for the case of $\kappa=1$, $t_{so}=0.5$, M=4M=-4, and $\phi=0$, under periodic boundary conditions (PBCs) for various values of $\lambda$. The fractal dimension of $n$th eigenstate in real space is defined as $D_{n}=-\lim_{L\to\infty}\ln{(\mathrm{IPR}_{n})}/\ln{(2L)}$, where the inverse participation ratio (IPR) is given by $\mathrm{IPR}_{n}=\sum_{i,\beta}|\psi_{i\beta}^{(n)}|^{4}$. In the thermodynamic limit, $D_{n}\to 0$ for localized states and $D_{n}\to 1$ for extended states. When $0<D_{n}<1$, the eigenstate is considered multifractal. Multifractality can also be detected by analyzing the fractal dimension in momentum space, defined as $D^{(k)}_{n}=-\lim_{L\to\infty}\ln{(\mathrm{IPR}^{(k)}_{n})}/\ln{(2L)}$, where $\mathrm{IPR}^{(k)}_{n}=\sum_{l}n^{2}_{k_{l}}$ with $k_{l}=2\pi l/L$ (l=1,,L)(l=1,...,L). For extended (localized) states in the real space, $D^{(k)}_{n}$ extrapolates to 0 $(1)$ , whereas in the multifractal regime, $D^{(k)}_{n}$ takes values significantly different from both 0 and $1$. As shown in Fig. 4, for both topologically trivial and nontrivial phases, the fractal dimensions of nearly all states in both real and momentum spaces remain system-size independent, and deviate from the limiting values of 0 and $1$. To further explore the thermodynamic limit, we analyze the scaling behavior of the mean fractal dimension in the real space, $\langle D_{r}\rangle=1/(2L)\sum_{n}D_{n}$, and in the momentum space, $\langle D_{k}\rangle=1/(2L)\sum_{n}D^{(k)}_{n}$, as shown in Fig. 5. In the thermodynamic limit, both $\langle D_{r}\rangle$ and $\langle D_{k}\rangle$ approach values distinct from both 0 and $1$ in all phases. Specially, we fit the scaling behavior for the both mean fractal dimensions: $\langle D_{r}\rangle=a_{r}/\ln(2L)+\langle D_{r}^{\infty}\rangle$ and $\langle D_{k}\rangle=a_{k}/\ln(2L)+\langle D_{k}^{\infty}\rangle$, respectively. Table 1 summarizes the extrapolated mean fractal dimensions $\langle D_{r}^{\infty}\rangle$ and $\langle D_{k}^{\infty}\rangle$ for selected values of $\lambda$ in different topological phases, with $\kappa=1$, $t_{so}=0.5$, M=4M=-4, and $\phi=0$. Our results indicate that eigenstates of the Fibonacci modulated chain exhibit robust multifractal characteristics across different topological phases, and the emergence of the TAI phase does not alter the underlying localization properties of the system.
Table 1: The extrapolated mean fractal dimensions $\langle D_{r}^{\infty}\rangle$ and $\langle D_{k}^{\infty}\rangle$ in different topological phases.
$\lambda$  2.0  3.0  3.5  3.8  4.2  4.8
$\langle D_{r}^{\infty}\rangle$  0.5850  0.5621  0.6727  0.5054  0.6667  0.5513
$\langle D_{k}^{\infty}\rangle$  0.5144  0.5857  0.4522  0.6912  0.4589  0.6144
  • a
    The system is localized in both topologically trivial (the quasiperiodic modulation amplitude $\lambda=2.0$, $3.5$, and $4.2$) and nontrivial ($\lambda=3.0$, $3.8$, and $4.8$) phases. Here, the spin-flip hopping strength $t_{so}=0.5$, the uniform Zeeman potential M=4M=-4, $\kappa=1$, and $\phi=0$.
Dynamical detection
Experimentally, this Fibonacci-modulated SOC chain can be realized along the momentum lattice Gadway2015 ; Gadway2016 ; Gadway2017 ; Gadway2018 ; Gadway201802 ; WangYF2022 ; LiYQ2022 ; YanBo2022 ; YanBo202202 ; LiYQ2023 ; YanBo2024 ; Gadway2024 in a cold atom of Bose-Einstein condensate, as illustrated in Fig. 6. In this setup, the spin and lattice-site degrees of freedom are encoded in the two ground-state hyperfine manifolds of selected atoms, such as 87Rb YanBo2024 . Specifically, the hyperfine state $F=1$ represents spin-down, while $F=2$ represents spin-up. A pair of counter-propagating laser beams is employed: the incoming laser beam has a fixed frequency, while the acousto-optic modulator in the reflected laser beam generate multi-frequency components. These laser beams drive a series of two-photon Bragg transitions between adjacent discrete momentum states. Transitions within the same hyperfine manifold enable nearest-neighbor couplings with the same spin, corresponding to the diagonal terms of $\mathcal{R}$. In contrast, transitions between different hyperfine manifolds facilitate nearest-neighbor couplings with opposite spins, corresponding to the off-diagonal terms of $\mathcal{R}$. By precisely adjusting the amplitude of each multi-frequency component, all couplings can be individually controlled. Additionally, the on-site potentials $\mathcal{U}_{i}$ can be tuned via the detuning of the two-photon Bragg transitions.
Figure 6: Illustration of experimental scheme. (a) A quasi-1D BEC is illuminated by a pair of counterpropagating laser beams: one with a fixed frequency $\omega$ and the other containing multifrequency components $\omega_{j}$. (b) The lasers, far detuned from the atomic transition, drive a series of engineered two-photon Bragg transitions. These transitions couple different momentum states within the same ground-state hyperfine manifold (blue and yellow arrows) or across different ground-state hyperfine manifolds (black arrows), with an increment of $2\hbar\tilde{k}$, where $\tilde{k}=2\pi/\tilde{\lambda}$ with $\tilde{\lambda}$ being the wave length of the lasers. The solid and dashed arrows represent the processes of photon absorption and emission, respectively.
Figure 7: Dynamical detections of topological and localization properties. (a) The disorder-averaged expectation value of the chiral displacement operator $\overline{C}(t)$ versus time $t$ for different quasiperiodic modulation amplitudes $\lambda$ with $N_{c}=50$ disorder realizations. (b) The time average of the chiral displacement operator $\langle\overline{C}\rangle$ as a function of the quasiperiodic modulation amplitude $\lambda$ with $N_{c}=200$ disorder realizations. The pink area represents that the system is localized in topologically nontrivial phase. (c) The evolution of the dynamical free energy $\Lambda(t)$ with $\lambda_{f}=0$, calculated for the $(L+1)$th eigenstate at different values of $\lambda_{i}$ as the initial states. (d) The evolution of the mean-square displacement $W(t)$ for different $\lambda$ with $\phi=0$. The black dashed line indicates the linear fit, used to quantify the slope magnitude. Here, the spin-flip hopping strength $t_{so}=0.5$, the uniform Zeeman potential M=4M=-4, and $L=144$.
@@
Topological properties and localization features of the system can be detected by the quench dynamics. To characterize the topology of our model, we monitor the dynamical response of the system to a sudden quench. Specifically, we measure the mean chiral displacement Gadway2018 ; Luzp2025 ; XueQK2024 ; Massignan2017 ; ZhangDW202202 and the Loschmidt echo (LE) ZhouBZ2021 ; ZhouBZ2019 ; Nehra2024 ; Langari2021 ; ChenB2024 ; HuZX2024 . We define the disorder-averaged expectation value of the chiral displacement operator as $\overline{C}(t)=2\langle\Psi(t)\left|\Gamma X\right|\Psi(t)\rangle$, where $\Gamma$ is the chiral operator, $X$ is the unit cell operator, and $|\Psi(t)\rangle=e^{iHt}|\Psi(0)\rangle$ is the time-evolved wave function, with the initial state $|\Psi(0)\rangle=c_{j_{0}\uparrow}^{\dagger}|0\rangle$ and $j_{0}$ being the position of the central bulk lattice site. The dynamics of $\overline{C}(t)$ generally exhibit transient, oscillatory behavior, as shown in Fig. 7(a) for different $\lambda$ with M=4M=-4 and $t_{so}=0.5$. In long time limit, $\bar{C}(t)$ converges to 0 for $\lambda$ values corresponding to the topologically trivial phase and to 1-1 for those corresponding to the nontrivial phase. To eliminate the oscillation, we compute the time average $\langle\overline{C}\rangle$, which converges to 1-1 for a topologically nontrivial phase and to 0 for a trivial one. Figure 7(b) shows $\langle\overline{C}\rangle$ as the function of the modulation amplitude $\lambda$ with M=4M=-4 and $t_{so}=0.5$. The value of $\langle\overline{C}\rangle$ starts at 0, then jumps to a nontrivial value three times, before eventually returning to 0. These results demonstrate that the mean chiral displacement is a sensitive probe of the multiple TAI phases in our model, effectively capturing the topological transitions driven by the modulation amplitude $\lambda$.
@@
The LE, defined as $\mathcal{L}(t)=|\langle\Psi(0)|e^{-iH(\lambda_{f})t}|\Psi(0)\rangle|^{2}$, is a powerful tool for analyzing nonequilibrium dynamics. It exhibits a series of zero points at specific time intervals when the initial Hamiltonian $H(\lambda_{i})$] and the post-quench Hamiltonian $H(\lambda_{f})$] describe systems that are localized in different phases. This behavior is a hallmark of dynamical quantum phase transitions (DQPTs), which have been observed in various systemsHeyl2013 ; Heyl2014 ; Heyl2015 ; Jafari2017 ; Heyl2018 ; LiJun2020 ; Bandyopadhyay2021 . Pioneering studies have established a connection between DQPTs and emergent topological transitions. In particular, dynamical topological order parameters can change their integer values when DQPTs occur, providing a means to dynamically track the topological changes during the quench. The appearance of zero points in the LE signals the nonanalytic behavior in the dynamical free energy $\Lambda(t)=-1/(2L)\ln\mathcal{L}(t)$. Figure 7(c) shows the evolution of $\Lambda(t)$ with $\lambda_{f}=0$, calculated for the $(L+1)$th eigenstate at different values of $\lambda_{i}$ for M=4M=-4 and $t_{so}=0.5$ as the initial states. For the topologically trivial Hamiltonian $H(\lambda_{f})$, when the initial Hamiltonian $H(\lambda_{i})$ is in the topologically nontrivial phase ($\lambda_{i}=3$, $3.8$, and $4.8$), a serial of divergence points of $\Lambda(t)$–corresponding to exact zeros of LE–emerge along the time axis. In contrast, for $\lambda_{i}$ values within the topologically trivial phase ($\lambda_{i}=2$, $3.5$, and $4.2$), the nonanalytic behavior disappears, indicating that both the initial and the post-quench Hamiltonians are localized in the same topological phase. These results suggest that the LE method not only captures the occurrence of DQPTs but also provides a powerful tool to probe the emergence of multiple TAI phases in our system.
@@
To dynamically detect the localization properties of our system, we employ the mean-square displacement Luzp2025 ; Rosch2012 ; BlochI2018 ; Wang2020 defined as
@@
Conclusion
We have demonstrated that multiple TAI phases can be induced by Fibonacci modulation in a SOC chain. Unlike conventional TAI phases, where the system exhibits full localization, our results show that the system retains multifractal features, regardless of whether it is in the TAI phase or not. Crucially, the emergence of multiple TAI phases is closely tied to the fractal structure of the spectrum induced by the Fibonacci modulation. The self-similar splitting of the bands at each scale enables the formation of multiple TAI phases, resulting in a topologically nontrivial structure that evolves as the SOC amplitude is varied. This model can be experimentally realized along the momentum lattice in a cold atomic system, and its properties can be tested through dynamical evolution experiments.
Methods
To determine the topological properties of the quasiperiodic-modulated SOC chain, we utilize the scattering matrix $S$, which relates the incoming and outgoing wave amplitudes at the Fermi level Zhang2016 :
$S=\begin{pmatrix}\tilde{R}_{\leftarrow}&\tilde{T}_{\leftarrow}\\ \tilde{T}_{\rightarrow}&\tilde{R}_{\rightarrow}\\ \end{pmatrix}$, (5)
[…]
where $\mathrm{sgn}[\cdots]$ denotes the sign function. A value of $Q=1$ indicates a topologically trivial phase, while Q=1Q=-1 signifies a topologically nontrivial phase. To account for the effects of disorder, we define the disorder-averaged topological invariant $\overline{Q}$, calculated over $N_{c}$ disorder realizations. In our analysis, $N_{c}$ configurations are sampled for different values of $\phi\in[0,2\pi]$.
$\begin{pmatrix}\mathcal{R}_{i+1,i}^{\dagger}\Psi_{i}\\ \Psi_{i+1}\end{pmatrix}=\tilde{B}_{i}\begin{pmatrix}\mathcal{R}_{i,i-1}^{\dagger}\Psi_{i-1}\\ \Psi_{i}\end{pmatrix}$ (7)
where
$\tilde{B}_{i}=\begin{pmatrix}0&\mathcal{R}_{i+1,i}^{\dagger}\\ -\mathcal{R}_{i,i+1}^{-1}&-\mathcal{R}_{i,i+1}^{-1}\mathcal{U}_{i}\\ \end{pmatrix}$. (8)
$B_{i}=U^{\dagger}\tilde{B}_{i}U,$ (9)
where
$U=\frac{1}{\sqrt{2}}\begin{pmatrix}I&I\\ iI&-iI\end{pmatrix}$, (10)
and $I$ is a $2\times 2$ identity matrix. In this base, the reflection (R~,R~)($\tilde$$_{\leftarrow},\tilde{R}_{\rightarrow}$) and transmission (T~,T~)($\tilde$$_{\leftarrow},\tilde{T}_{\rightarrow}$) matrices can be determined via the relations:
$\begin{pmatrix}\tilde{T}_{\rightarrow}\\ 0\end{pmatrix}=B\begin{pmatrix}I\\ \tilde{R}_{\leftarrow}\end{pmatrix},\quad\begin{pmatrix}\tilde{R}_{\rightarrow}\\ I\end{pmatrix}=B\begin{pmatrix}0\\ \tilde{T}_{\leftarrow}\end{pmatrix}$, (11)
where $B=B_{L}B_{L-1}\ldots B_{1}$.
Data Availability
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Code availability
The code used for the analysis is available from the authors upon reasonable request.
Acknowledgements